OpenStudy (anonymous):

Question on Ellipses! Please Help! Just check it for me, I' already answered it! Please?

OpenStudy (anonymous):

looks like length of major axis is 10, from -5 to 5 or am i mistaken?

OpenStudy (anonymous):

i think am right

OpenStudy (anonymous):

because the left hand endpoint of the ellipse is \((-5,0)\) and the right hand endpoint is \((5,0)\) the the distance between those two points is 10 by counting

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok hold on and let me make sure i did not tell you wron g

OpenStudy (anonymous):

no what i said was correct

OpenStudy (anonymous):

ok lets go a little slower the first answer was ten but the foci are not \((-5,0)\) and \((5,0)\)

OpenStudy (anonymous):

for the last problem your "co vertex" is correct

OpenStudy (anonymous):

for the first question you asked, the ellipse is given by the equation \[\frac{x^2}{25}+\frac{y^2}{9}=1\] and i know that because \(5^2=25\) and \(3^2=9\) and the center was at \((0,0)\)

OpenStudy (anonymous):

to find the foci, you need to compute \[c^2=a^2-b^2\] \[c^2=25-9\] \[c^2=16\] \[c=4\]

OpenStudy (anonymous):

since the center is at \((0,0)\) and the foci are 4 units to the right and left of the center, they are \((-4,0)\) and \((0,4)\)

OpenStudy (anonymous):

notice i did not really have to find the equation of the ellipse you see the 5 and the 3 from your eyes, so you just compute \[c^2=5^2-3^2\] to get the foci

OpenStudy (anonymous):

the first question your answer to the second question is correct

OpenStudy (anonymous):

i have plenty, and i can't even cash them in for valuable prizes, so no worries